OpenStudy (anonymous):

A water skier is pulled by a net force of 180 N and has an acceleration of 2 m/s2. What is the mass of the skier? A. 80 kg B. 360 kg C. 90 kg D. 60 kg

OpenStudy (shamim):

Required formula is F=ma

OpenStudy (shamim):

Given F=180N a=2m/s m=?

OpenStudy (shamim):

Feel free to ask for more clarification

OpenStudy (anonymous):

c f=ma 180=m*2, m=90